# EMMA: End-to-End Multimodal Model for Autonomous Driving

### v2

@@ Abstract
We introduce EMMA, an End-to-end Multimodal Model for Autonomous driving. Built on a multi-modal large language model foundation, EMMA directly maps raw camera sensor data into various driving-specific outputs, including planner trajectories, perception objects, and road graph elements. EMMA maximizes the utility of world knowledge from the pre-trained large language models, by representing all non-sensor inputs (e.g. navigation instructions and ego vehicle status) and outputs (e.g. trajectories and 3D locations) as natural language text. This approach allows EMMA to jointly process various driving tasks in a unified language space, and generate the outputs for each task using task-specific prompts. Empirically, we demonstrate EMMA’s effectiveness by achieving state-of-the-art performance in motion planning on nuScenes as well as competitive results on the Waymo Open Motion Dataset (WOMD). EMMA also yields competitive results for camera-primary 3D object detection on the Waymo Open Dataset (WOD). We show that co-training EMMA with planner trajectories, object detection, and road graph tasks yields improvements across all three domains, highlighting EMMA’s potential as a generalist model for autonomous driving applications. However, EMMA also exhibits certain limitations: it can process only a small amount of image frames, does not incorporate accurate 3D sensing modalities like LiDAR or radar and is computationally expensive. We hope that our results will inspire further research to mitigate these issues and to further evolve the state of the art in autonomous driving model architectures.
@@ 1 Introduction
Historically, autonomous driving systems employed a modular approach, consisting of specialized components for perception (Yurtsever et al., 2020; Li et al., 2022b; Lang et al., 2019; Sun et al., 2022; Hwang et al., 2022), mapping (Li et al., 2022a; Tancik et al., 2022), prediction (Nayakanti et al., 2023; Shi et al., 2024), and planning (Teng et al., 2023). While this design lends itself to easier debugging and optimization of individual modules, it poses scalability challenges due to the accumulated errors among modules and limited inter-module communication. In particular, the expert-designed interfaces between modules, such as the perception and behavior modules, may struggle to adapt to novel environments because they are often pre-defined based on targeted scenarios (Bansal et al., 2019; Jiang et al., 2023; Nayakanti et al., 2023; Seff et al., 2023). End-to-end autonomous driving systems (Hu et al., 2023; Zhai et al., 2023; Li et al., 2024) have recently emerged as a potential solution, directly learning to generate driving actions from sensor data. This approach eliminates the need for symbolic interfaces between modules and allows for joint optimization of driving objectives from raw sensor inputs. However, these systems are often specialized for specific driving tasks and trained on limited datasets, hindering their ability to generalize to rare or novel scenarios.
@@
Multimodal Large Language Models (MLLMs) (Gemini Team Google, 2023; Achiam et al., 2023) offer a promising new paradigm for AI in autonomous driving that may help to address such challenges. This is because MLLMs, as generalist foundation models, excel in two key areas: (1) they are trained on vast, internet-scale datasets that provide rich "world knowledge" beyond what is contained in common driving logs, and (2) they demonstrate superior reasoning capabilities through techniques such as chain-of-thought reasoning (Wei et al., 2022) that are not available in specialized driving systems. While recent efforts (Chen et al., 2024b; Tian et al., 2024) have explored integrating and augmenting the capabilities of existing driving systems with MLLMs, we propose to develop an autonomous driving system in which the MLLM is a first class citizen.
@@
We introduce the End-to-End Multimodal Model for Autonomous Driving (EMMA), built on top of Gemini (Gemini Team Google, 2023) without additional specialized components. Figure 1 shows the overview of the EMMA framework. EMMA  accepts camera images and plain text for other non-vision inputs such as high-level driving commands and historical context. By recasting driving tasks as visual question answering (VQA) problems, EMMA  leverages Gemini’s pre-trained capabilities while preserving its extensive world knowledge. After EMMA is fine-tuned with driving logs from all tasks using task-specific prompts (see Figure 2 for more examples), its shared Gemini model generates various driving outputs such as future trajectories for motion planning, perception objects, road graph elements, and scene semantics. Our experiments showcase EMMA’s strong performance on several planning and perception benchmarks despite this simple design. Additionally, we find that EMMA  can produce interpretable, human-readable outputs for many perception tasks such as road graph estimation, and is able to function as a generalist model that is both scalable and robust for autonomous driving. Notably, as used here and throughout the paper, the EMMA generalist model refers to a machine learning model that has been trained and fine-tuned on a large volume of driving data to perform a wide range of specific driving tasks in the autonomous driving domain.
@@
Despite these promising results, EMMA is not without its limitations. In particular, it faces challenges for real-world deployment due to: (1) limitations in 3D spatial reasoning due to its inability to fuse camera inputs with LiDAR or radar, (2) the need for realistic and computationally expensive sensor simulation to power its closed-loop evaluation, and (3) the increased computational requirements relative to conventional models. We plan to better understand and address such challenges in future work.
In the remainder of this paper, Section 2 describes the detailed method of EMMA for end-to-end motion planning and generalist tasks in autonomous driving. In Section 3, we present experimental results of EMMA on public and internal datasets. Finally, we discuss related works and limitations of our work in Sections 4 and 5, respectively.
@@ 2 Method
Figure 1: EMMA overview diagram. It takes 3 inputs (left): 1) a high-level command from the router, 2) historical status of the ego vehicle, and 3) surround-view camera videos. The model then predicts ego future trajectories (right) for motion planning that will be transformed into vehicle driving control signals. Further, we can ask the model to explain its rationale (top right) before predicting trajectories, which enhances both the performance and explainability of the model through chain-of-thought reasoning. Notably, we incorporate visual grounding into the rationale so that the model also predicts the accurate 3D/BEV location for critical objects. For completeness, we highlight three additional capabilities of our model (bottom).
EMMA is built on top of Gemini (Gemini Team Google, 2023), a family of MLLMs developed at Google. We leverage the auto-regressive Gemini models that are trained to process interleaved textual and visual inputs to produce text outputs:
@@
As shown in Figure 1, we map autonomous driving tasks into Gemini-based EMMA formulation. All sensor data are represented as stitched images or videos as $\mathbf{V}$; all router commands, driving context, and task-specific prompts are represented as $\mathbf{T}$; all output tasks are presented as language outputs $\mathbf{O}$. A challenge is that many of the inputs and outputs need to capture 3D world coordinates, such as waypoint BEV (Bird’s Eye View) locations $(x,y)$ for motion planning and the location and size of 3D boxes. We consider two representations: The first is direct text conversion to floating-point numbers, expressed as $\mathbf{T_{\text{coordinates}}}=\{(x_{i},y_{i})\}\approx\texttt{text}(\{(x_{i},y_{i})\})$, where the specified decimal precision depends on the distance unit and decimal points. RT-2 (Brohan et al., 2023) exemplifies this approach in robotic control. The second approach uses special tokens to represent each location or action, formulated as $\mathbf{T_{\text{coordinates}}}=\{(x_{i},y_{i})\}\approx\texttt{tokenize}(\{(x_{i},y_{i})\})$, with resolution determined by the learned or manually defined discretization scheme. MotionLM (Seff et al., 2023) leverages this method for motion forecasting. We note that the two approaches have their respective strengths and weaknesses. We opt for the text representation such that all tasks can share the same unified language representation space and they can maximally reuse the knowledge from the pre-trained weights, even though the text presentation may produce more tokens than specialized tokenization.
@@ 2.1 End-to-End Motion Planning
While we are not the first to adopt this general formulation—(Li et al., 2024) conducted a thorough investigation, particularly examining the impact of including the historical ego status—our contribution lies in adapting this formulation specifically for MLLMs for autonomous driving.
@@ 1. R1 -
    Scene description broadly describes the driving scenarios, including weather, day of time, traffic situations, and road conditions. For example: The weather is clear and sunny, and it is daytime. The road is four-lane undivided street with a crosswalk in the middle. There are cars parked on both sides of the street.
@@ 3. R3 -
    Behavior description of critical objects describes the current status and intent for the identified critical objects. A concrete example is as follows: The pedestrian is currently standing on the sidewalk, looking toward the road, and maybe preparing to cross the street. The vehicle is currently ahead of me, moving in the same direction, and its future trajectory suggests it will continue straight.
@@ 4. R4 -
We highlight that the driving rationale captions are generated using an automated tool without any additional human labels, ensuring scalability of the data generation pipeline. Specifically, we leverage off-the-shelf perception and prediction expert models to identify critical agents, and then use Gemini models with carefully designed visual and text prompts to generate comprehensive scene and agent behavior descriptions. Meta driving decisions are computed using a heuristic algorithm that analyzes the ego vehicle’s ground-truth trajectory.
@@ 2.3 EMMA Generalist
Our vision-language framework represents all non-sensor inputs and outputs as plain text, providing the flexibility necessary to incorporate many other driving tasks. We employ instruction-tuning, a well-established approach in LLMs, to jointly train all tasks together with task-specific prompts included in the inputs $\mathbf{T}$ of Eq. 1. We organize these tasks into three primary categories: spatial reasoning, road graph estimation, and scene understanding. Fig. 2 illustrates the overall EMMA Generalist diagram.
@@
We then represent the detection tasks using a fixed prompt $\mathbf{T_{\text{detect\_3D}}}$, such as “detect every object in 3D”, as in the following:
@@
While $\mathbf{O_{\text{boxes}}}$ is an unordered set of boxes, the predictions from an auto-regressive language model are always ordered. We find that sorting the 3D bounding boxes by depth improves detection quality, in contrast with the findings in Pix2Seq (Chen et al., 2022a).
@@
While edges within each polyline are directional, each polyline does not necessarily have a unique order relative to the other elements. This bears similarity to object detection (e.g., (Carion et al., 2020; Chen et al., 2022a)), where each box is defined by ordered attributes (top-left corner, bottom-right corner), but a relative ordering between boxes does not necessarily exist. There are several existing works that model polyline graphs with Transformers (Yuan et al., 2024; Liao et al., 2024a, b, 2023; Ding et al., 2023; Qiao et al., 2023; Liu et al., 2023; Li et al., 2022a), sharing similarities with language models.
@@ 2.4 Generalist Training
Our unified vision-language formulation enables the simultaneous training of multiple tasks with a single model, allowing for task-specific predictions at inference time through simple variations of the task prompt $\mathbf{T}_{\text{task}}$. The training procedure is both straightforward and flexible.
@@ 3 Experiments
We highlight the experiments used to validate the efficacy of EMMA model. We leverage the smallest size of Gemini, i.e., Gemini 1.0 Nano-1 (Gemini Team Google, 2023) for all the experiments. In the following section, we first present the results of end-to-end planner trajectory generation on two public datasets in Section 3.1. Next, we conduct additional experiments on our internal datasets, studying the impact of chain-of-thought and data scaling in Section 3.2. Section 3.3 focuses on experiments involving three core perception tasks: 3D object detection, road graph estimation, and scene understanding. Our co-training results for the generalist model are summarized in Section 3.4. Finally, we showcase visual results that highlight EMMA capabilities in challenging, long-tail scenarios in Section 3.5.

### v3

@@ Abstract
We introduce EMMA, an End-to-end Multimodal Model for Autonomous driving. Built upon a multi-modal large language model foundation like Gemini, EMMA directly maps raw camera sensor data into various driving-specific outputs, including planner trajectories, perception objects, and road graph elements. EMMA maximizes the utility of world knowledge from the pre-trained large language models, by representing all non-sensor inputs (e.g. navigation instructions and ego vehicle status) and outputs (e.g. trajectories and 3D locations) as natural language text. This approach allows EMMA to jointly process various driving tasks in a unified language space, and generate the outputs for each task using task-specific prompts. Empirically, we demonstrate EMMA’s effectiveness by achieving state-of-the-art performance in motion planning on nuScenes as well as competitive results on the Waymo Open Motion Dataset (WOMD). EMMA also yields competitive results for camera-primary 3D object detection on the Waymo Open Dataset (WOD). We show that co-training EMMA with planner trajectories, object detection, and road graph tasks yields improvements across all three domains, highlighting EMMA’s potential as a generalist model for autonomous driving applications. We hope that our results will inspire research to further evolve the state of the art in autonomous driving model architectures.
@@ 1 Introduction
Historically, autonomous driving systems employed a modular approach, consisting of specialized components for perception (Yurtsever et al., 2020; Li et al., 2022b; Lang et al., 2019; Sun et al., 2022; Hwang et al., 2022), mapping (Li et al., 2022a; Tancik et al., 2022), prediction (Nayakanti et al., 2023; Shi et al., 2024), and planning (Teng et al., 2023; Lioutas et al., 2022). While this design lends itself to easier debugging and optimization of individual modules, it poses scalability challenges due to the limited inter-module communication. In particular, the expert-designed interfaces between modules, such as the perception and behavior modules, may struggle to adapt to novel environments because they are often pre-defined based on targeted scenarios (Bansal et al., 2019; Jiang et al., 2023; Nayakanti et al., 2023; Seff et al., 2023). End-to-end autonomous driving systems (Hu et al., 2023; Zhai et al., 2023; Li et al., 2024) have recently emerged as a potential solution, directly learning to generate driving actions from sensor data. This approach eliminates the need for symbolic interfaces between modules and allows for joint optimization of driving objectives from raw sensor inputs. However, these systems are often specialized for specific driving tasks and trained on limited datasets, hindering their ability to generalize to rare or novel scenarios.
@@
Multimodal Large Language Models (MLLMs) (Gemini Team Google, 2023; Achiam et al., 2023) offer a promising new paradigm for AI in autonomous driving that may help to address such challenges. This is because MLLMs, as generalist foundation models, excel in two key areas: (1) they are trained on vast, internet-scale datasets that provide rich "world knowledge" beyond what is contained in common driving logs, and (2) they demonstrate superior reasoning capabilities through techniques such as chain-of-thought reasoning (Wei et al., 2022; Zhang et al., 2023b) that are not available in specialized driving systems. While recent efforts (Chen et al., 2024b; Tian et al., 2024) have explored integrating and augmenting the capabilities of existing driving systems with MLLMs, we propose to develop an autonomous driving system in which the MLLM is a first class citizen.
@@
We introduce the End-to-End Multimodal Model for Autonomous Driving (EMMA), built on top of a multimodal large language model, such as Gemini (Gemini Team Google, 2023) or PaLI (Chen et al., 2024d) without additional specialized components. Figure 1 shows the overview of the EMMA framework. EMMA accepts camera images and plain text for other non-vision inputs such as high-level driving commands and historical context. By recasting driving tasks as visual question answering (VQA) problems, EMMA leverages Gemini’s pre-trained capabilities and extensive world knowledge. After EMMA is fine-tuned with driving logs from all tasks using task-specific prompts (see Figure 2 for more examples), it generates various driving outputs such as future trajectories for motion planning, perception objects, road graph elements, and scene semantics. Our experiments showcase EMMA’s strong performance on several planning and perception benchmarks despite this simple design. Additionally, we find that EMMA can produce interpretable, human-readable outputs for many perception tasks such as road graph estimation, and is able to function as a generalist model that is both scalable and robust for autonomous driving. Notably, as used here and throughout the paper, the EMMA generalist model refers to a machine learning model that has been trained and fine-tuned on a large volume of driving data to perform a wide range of driving tasks in the autonomous driving domain.
@@
In the remainder of this paper, Section 2 describes the detailed method of EMMA for end-to-end motion planning and generalist tasks in autonomous driving. In Section 3, we present experimental results of EMMA on public and internal datasets. Finally, we discuss related works in Sections 4.
Despite these promising results, EMMA is not without its limitations. We discuss the limitations in-depth in the Appendix Section A.5. In particular, it faces challenges for real-world deployment due to: (1) limitations in 3D spatial reasoning due to its inability to fuse camera inputs with LiDAR or radar, (2) the need for realistic and computationally expensive sensor simulation to power its closed-loop evaluation, and (3) the increased computational requirements relative to conventional models. We plan to better understand and address such challenges in future work.
@@ 2 Method
Figure 1: EMMA overview diagram. It takes 3 inputs (left): 1) a high-level command from the router, 2) historical status of the ego vehicle, and 3) surround-view camera videos. The model then predicts ego future trajectories (right) for motion planning that will be transformed into vehicle driving control signals. Further, we can ask the model to explain its rationale (top) before predicting trajectories, which enhances both the performance and explainability of the model through chain-of-thought reasoning. Notably, we incorporate visual grounding into the rationale so that the model also predicts the accurate 3D/BEV location for critical objects. In addition to end-to-end planning, we highlight three additional perception capabilities of our model (bottom).
While we will show EMMA can be compatible with various MLLMs such as Gemini (Gemini Team Google, 2023) and PaLI (Chen et al., 2024d) in our experiments, this section will focus on our main EMMA based on Gemini. We leverage the auto-regressive Gemini models that are trained to process interleaved textual and visual inputs to produce text outputs:
@@
As shown in Figure 1, we map autonomous driving tasks into our Gemini-based EMMA formulation. All sensor data are represented as stitched images or videos $\mathbf{V}$; all router commands, driving context, and task-specific prompts are represented in language prompts $\mathbf{T}$; and all output tasks are presented as language outputs $\mathbf{O}$. A challenge is that many of the inputs and outputs need to capture 3D world coordinates, such as waypoint BEV (Bird’s Eye View) locations $(x,y)$ for motion planning and the location and size of 3D boxes. We consider two representations: The first is direct text conversion to floating-point numbers, expressed as $\mathbf{T_{\text{coordinates}}}=\{(x_{i},y_{i})\}\approx\texttt{text}(\{(x_{i},y_{i})\})$, where the specified decimal precision depends on the distance unit and decimal points. RT-2 (Brohan et al., 2023) exemplifies this approach in robotic control. The second approach uses special tokens to represent each location or action, formulated as $\mathbf{T_{\text{coordinates}}}=\{(x_{i},y_{i})\}\approx\texttt{tokenize}(\{(x_{i},y_{i})\})$, with resolution determined by the learned or manually defined discretization scheme. MotionLM (Seff et al., 2023) leverages this method for motion forecasting. We note that the two approaches have their respective strengths and weaknesses. We opt for the text representation such that all tasks can share the same unified language representation space and they can maximally reuse the knowledge from the pre-trained weights, even though the text representation may produce more tokens than specialized tokenization.
@@ 2.1 End-to-End Motion Planning
While we are not the first to adopt this general formulation—(Li et al., 2024) conducted a thorough investigation, particularly examining the impact of including the historical ego status—our contribution lies in adapting this formulation specifically for MLLMs for autonomous driving. Our self-supervised approach exists alongside other notable methods that explore reconstruction via spatio-temporal scene decomposition (Yang et al., 2023), world modeling (Zhang et al., 2023a), or joint motion prediction (Wagner et al., 2024). Beyond this self-supervised foundation, the following sections explore enhancements for EMMA by incorporating reasoning and developing generalist setups with human or auto-labels.
@@ 1. R1 -
    Scene description broadly describes the driving scenarios, including weather, day of time, traffic situations, and road conditions. We provide a concrete example for prompting Gemini. Example prompt: Assume you are an autonomous vehicle, and the images come from your front cameras. Can you describe the current scenario in terms of weather, time of the day, road environment, lane options, and your ego lane position? Example answer: The weather is clear and sunny, and it is daytime. The road is four-lane undivided street with a crosswalk in the middle. There are cars parked on both sides of the street.
@@ 3. R3 -
    Behavior description of critical objects describes the current status and intent for the identified critical objects. We provide a concrete example for prompting Gemini. Example prompt: Assess the potential risks posed by the [focused_agent] with red bounding box. Summarize any immediate concerns that need addressing to maintain safety, paying close attention to how the objects may affect your route. Example answer: The pedestrian is currently standing on the sidewalk, looking toward the road, and maybe preparing to cross the street. The vehicle is currently ahead of me, moving in the same direction, and its future trajectory suggests it will continue straight.
@@ 4. R4 -
We highlight that the driving rationale captions are generated using an automated tool without any additional human labels, ensuring scalability of the data generation pipeline. Specifically, we leverage off-the-shelf perception and prediction expert models to identify critical agents, and then use Gemini models with carefully designed visual and text prompts to generate scene and agent behavior descriptions. Meta driving decisions are computed using a heuristic algorithm that analyzes the ego vehicle’s ground-truth trajectory.
@@ 2.3 EMMA Generalist
Our vision-language framework represents all non-sensor inputs and outputs as plain text, providing the flexibility necessary to incorporate many other driving tasks. We employ instruction-tuning, a well-established approach in LLMs, to jointly train all tasks together with task-specific prompts included in the inputs $\mathbf{T}$ of Eq. 1. We organize these tasks into three primary categories: spatial reasoning, road graph estimation, and scene understanding.
@@
We then represent the detection tasks using a fixed prompt $\mathbf{T_{\text{detect\_3D}}}$, such as “detect every object in 3D”, as follows:
@@
While $\mathbf{O_{\text{boxes}}}$ is an unordered set of boxes, the predictions from an auto-regressive language model are always ordered. We find that sorting the 3D bounding boxes by depth improves detection quality, unlike the findings in Pix2Seq (Chen et al., 2022a).
@@
While edges within each polyline are directional, each polyline does not necessarily have a unique order relative to the other elements. This bears similarity to object detection (e.g., (Carion et al., 2020; Chen et al., 2022a)), where each box is defined by ordered attributes (top-left corner, bottom-right corner), but a relative ordering between boxes does not necessarily exist. There are several existing works that model polyline graphs with Transformers (Yuan et al., 2024; Liao et al., 2024a; b; 2023; Ding et al., 2023; Qiao et al., 2023; Liu et al., 2023; Li et al., 2022a), sharing similarities with language models.
@@ 2.4 Generalist Training
Our unified vision-language formulation enables the simultaneous training of multiple tasks with a single model, allowing for task-specific predictions at inference time through simple variations of the task prompt $\mathbf{T}_{\text{task}}$. This training procedure is both straightforward and flexible.
@@ 3 Experiments
Our experiments are primarily based on Gemini 1.0 Nano-1 (Gemini Team Google, 2023), with additional results provided for a variant of EMMA based on PaLI (Chen et al., 2024d). We first summarize the main datasets used for various experiments in Section 3.1. And then we present the results of end-to-end planner trajectory generation on two public datasets in Section 3.2. Next, we conduct experiments on our internal datasets, studying the impact of chain-of-thought and data scaling in Section 3.3. Section 3.4 focuses on 3D object detection experiments. Our co-training results for the generalist model are summarized in Section 3.5. Finally, we showcase visual results that highlight EMMA’s capabilities in challenging, long-tail scenarios in Section 3.6.
3.1 Summary of Datasets
Before diving into the experimental details on how we validate EMMA, we summarize the main datasets in this section. Overall, we leverage three public datasets, nuScenes (Caesar et al., 2020), Waymo Open Motion Dataset (WOMD) (Chen et al., 2024a) and Waymo Open Dataset (WOD) (Sun et al., 2020). We also constructed three large-scale internal datasets for end-to-end motion planning, 3D detection and road graph estimation. We summarize the dataset sizes in Table 1.
The nuScenes dataset (Caesar et al., 2020) offers a comprehensive autonomous vehicle sensor suite for evaluation. It consists of 1,000 scenes, each spanning 20 seconds, and includes information from 6 cameras that collectively provide 360-degree coverage in the field of view.
The WOMD dataset comprises 103k real-world urban and suburban driving scenarios, each lasting 20 seconds. These scenarios are further segmented into 1.1M examples, each representing a 9-second window: 1 second is used as input context, and the remaining 8 seconds serve as the prediction target. The dataset includes detailed map features such as traffic signal states and lane characteristics, along with agent states such as position, velocity, acceleration, and bounding boxes.
We build a large-scale internal motion planning dataset, boasting over 24 million real-world driving scenarios, each 30 seconds long. This makes it roughly 355 times larger than WOMD or any other publicly available driving dataset. To efficiently leverage this massive scale, we sample just one frame per scenario, yielding 24 million diverse training examples. This approach maximizes dataset diversity while maintaining computational efficiency.
We also construct two separate, large-scale internal datasets for detection and road graph tasks, comprising 12 million and 8 million examples, respectively. For the detection dataset, we prioritized scenarios with diverse objects, sampling one example every 3 seconds from these scenarios. The road graph dataset, on the other hand, focuses on diverse scenarios and geo-locations, so we sample one example every 30 seconds, aligning with our motion-planning dataset.
Lastly, we also validate the camera-based 3D object detection task on the public WOD benchmark (Sun et al., 2020; Hung et al., 2024). This benchmark offers 1150 20-second scenes, each providing meticulously synchronized and calibrated high-quality LiDAR, camera, and 3D box data from a variety of urban and suburban environments.
Dataset Name Total Hours of Driving Number of Training Examples
nuScenes (Caesar et al., 2020) 66 $18,686$
WOMD (Chen et al., 2024a) $572$ $487,061$
Internal Motion Planning Dataset 203,117203,117(355x) 24,374,04624,374,046(50x)
WOD (Sun et al., 2020) 66 $158,081$
Internal Detection Dataset $6250$ $11,765,140$
Internal Roadgraph Dataset $64,135$ $8,304,671$
Table 1: Summary of main training dataset scales. This table details the scales of the three public datasets (nuScenes, WOMD, WOD) and three large-scale internal datasets leveraged for studying data scaling and generalist properties.